A national park is famous for its beautiful desert landscape and its many natural rock formations. The following table is based on information gathered by a park ranger of all rock formations of at least 3 feet.  The height of the rock formation is rounded to the nearest foot.

Height of rock formation, feet 3
Answer #1

Total number of peaks = 117 + 89 + 32 + 31 + 5 = 274


(a) P(3 to 9 feet) = 117/274 = 0.427

(b) P(30 feet or taller) = (32+31+5)/274 = 0.248

(c) P(3 to 49 feet) = (117 + 89 + 32)/274 = 0.869

(d) P(10 to 74 feet) = (89+32+31)/274 = 0.555

(e) P( 75 feet or taller) = 5/274 = 0.018
Answer #2

117/(117+89+32+31+5)= 117/274=0.427

68/274= 0.248

238/274=0.869

152/274= 0.555

5/274= 0.018
Answer #3

A) 111 96 24 26 8=265. 111/265=0.42 (you round it)
B) 24 26 8=58. 58/265=0.22
C)111 96 24=231. 231/265=0.87
D)96 24 26=146. 146/265=0.55
E)8/265=0.03